Find the roots of the quadratic equation:
z^2 + 5z - 14 = 0

Answer by checkley71(8403)

Z~2+5Z-14=0 OR (Z+7)(Z-2)=0 OR Z+7=0 OR Z=-7 & Z-2=0 OR Z=2